Use the normal distribution of fish lengths for which the mean is 9 inches and the standard deviation is 4 inches. Assume the variable x is normally distributed.

What percentage of the fish are longer than 13 inches?
Answer #1

Solution :

Given ,

mean = = 9

standard deviation = = 4

P(x >13 ) = 1 - P(x< 13)

= 1 - P[(x -) / < (13-9) / 4]

= 1 - P(z <1 )

Using z table

= 1 - 0.5

= 0.5

answer=50%